OpenStudy (anonymous):

what is the square root of400 over 5?

OpenStudy (watchmath):

\[\sqrt{400}/5\] like that?

OpenStudy (anonymous):

If it's like watchmath's = 4 If not = \[4\sqrt{5}\]

OpenStudy (anonymous):

\[4\sqrt{5} \]